OpenStudy (anonymous):

Solve. 7^x – 5 = 7^4x + 7 A)x = –4 B) x = -12/5 C)x = 2/3 D) x = 4 x = negative 12 over 5 x = 2 over 3 x = 4

OpenStudy (nurali):

7x-5=4x+7 7x-4x = 7+5 =12 3x =12 x =12/3 = 4

OpenStudy (whpalmer4):

Is that supposed to be \[7^{x-5} = 7^{4x+7}\]? You really need to take the order of evaluation into account if you aren't going to write your problems with the equation editor. What you wrote is\[7^x-5 = 7^{4x}+7\]which is an entirely different problem.

OpenStudy (whpalmer4):

If the problem is as I asked, then you can simply solve \[x-5=4x+7\] to get your answer, because the bases on both sides are equal.

OpenStudy (whpalmer4):

@Nurali's solution is for an entirely different problem.